# Project Page: https://yikai-wang.github.io/asuka/ MISATO Dataset: https://github.com/Yikai-Wang/asuka-misato

### v2

@@ 4 Evaluation Dataset
Combined together, MISATO contains 2000 examples from four inpainting domains, indoor, outdoor landscape, building, and background. As shown in Fig. 6(b), we adopt the masking strategy as in Sec. 3.1 but excluding the rectangle and complement rectangle masks. The masking ratio is set as $[0.2,0.8]$.
5 Experiments
Implementation details. ASUKA utilizes the stable diffusion text-guided inpainting model v1.5 as the generative backbone. We train ASUKA on Places2 [61]. For the MAE [19] used in ASUKA, we train on images of size $256^{2}$, which is efficient and produce context-stable guidance for SD to generate high-resolution images. We fine-tune from the pre-trained MAE with a batch size of 1024. We use four self-attention blocks for the alignment module, training from scratch with AdamW [31] of learning rate 5e-2 with the standard diffusion objective. We set $p$ as $100\%$ and linearly decay it to $10\%$ in the first 2K training steps and then freeze. We fine-tune the decoder based on [62] for 50K steps with a batch size of 40 and learning rate of 8e-5 with cosine decay. We use ColorJitter for color augmentation, with brightness 0.15, contrast 0.2, saturation 0.1, and hue 0.03.
Evaluation metrics. We use the Learned Perceptual Image Patch Similarity (LPIPS) [58] to calculate the patch-level image distances, Fréchet Inception Distance (FID) [20] to compare the distribution distance between generated images and real images, and Paired/Unpaired Inception Discriminative Score (P-IDS/U-IDS) [59] to measure the human-inspired linear separability.
Competitors. We consider three variants of SD [39]: i) SD, prompted by "", performing unconditional generation; ii) SD-text, prompted by "background" as we have no captions in inpainting; iii) SD-prompt, prompted by learnable tokens which are trained using the same training pipeline of ASUKA. We also compare with state-of-the-art inpainting algorithms Co-Mod [59], MAT [25], LaMa [45], MAE-FAR [7], and SD-Repaint [32].
Table 1: Quantitative comparison on MISATO and Places 2. Co-Mod [59], MAT [25], LaMa [45], MAE-FAR [7] and SD-Repaint [32] are state-of-the-art inpainting methods. SD [39] performs unconditional generation. SD-text uses “background” text prompt to guide generation. SD-prompt uses learnable prompts trained specifically for inpainting, using the same training setting as ASUKA, performing prompt-guided generation. ASUKA and SD variants use the stable diffusion text-guided inpainting model v1.5.
Dataset MISATO (2k images) Places 2 (36.5k images)
@@ 5 Experiments
Co-Mod [59] 0.179 17.421 0.243 0.109 0.267 5.794 0.274 0.096
MAT [25] 0.176 17.261 0.255 0.122 0.202 3.765 0.348 0.195
LaMa [45] 0.155 15.436 0.260 0.135 0.202 6.693 0.247 0.050
MAE-FAR [7] 0.142 13.283 0.282 0.153 0.174 3.559 0.307 0.105
SD [39] 0.168 12.812 0.345 0.211 0.193 1.514 0.375 0.207
SD-text 0.164 12.603 0.337 0.207 0.191 1.506 0.373 0.202
SD-prompt 0.160 12.517 0.331 0.204 0.189 1.477 0.390 0.234
SD-Repaint [32] 0.227 27.861 0.016 0.007 0.251 12.466 0.217 0.045
ASUKA 0.150 11.460 0.368 0.256 0.183 1.230 0.413 0.287
5.1 Comparison on Benchmarks
Quantitative comparison results are reported in Tab. 1. ASUKA, although based on frozen SD, achieves significant and consistent improvements over SD across all the evaluation metrics, and enjoys the state-of-the-art results on FID, U-IDS and P-IDS. Particularly, the U-IDS and P-IDS coincide with human preference [59], and has a potential upper bound of $0.5$, indicating the superiority of ASUKA by a large margin. The inconsistency results between LPIPS and FID, U-IDS, P-IDS, partially comes from the frozen SD, where ASUKA achieves consistent improvements but is bounded to the frozen U-Net. In summary, the results validate the contribution of ASUKA that yields context-stable and visual-consistent inpainting through frozen SD. This improvement is both validated in in-distribution dataset Places 2 and out-of-distribution dataset MISATO.
Qualitative comparison examples are shown in Fig. 8. i) The state-of-the-art inpainting algorithms usually suffer from unnatural generation, for example the unnatural boundaries in the third and fourth rows, and failed inpainting of tower in the third-to-last row. LaMa and MAE-FAR sometimes lead to blurred inpainting results, especially in the scenario of large continuous masks. ii) The SD variants usually suffer from the context-stability issue and hallucinate unreasonable objects, in almost all the illustrated images. iii) Particularly, for the background mask shown in the second-to-last row, state-of-the-art inpainting algorithms cannot identify the object precisely, and extend the human in hairs and elbows. SD hallucinates random elements in the masked region. iv) In contrast, ASUKA enjoys context-stable and visual-consistent inpainting.
5.2 Further Analysis of ASUKA
Ablation of MAE prior. We compare our fine-tuned MAE with directly adopting the MAE trained in [7]. To this end, we train ASUKA with the MAE in [7] using the same training strategy and compare the results in Tab. 2(a). Results suggest the improvements of fine-tuning MAE, especially on FID and U-IDS. This improvement comes from the better adaptation on the real-world masks.
Ablation of alignment module. Our alignment module includes a dimension and a distribution alignment. We validate the effectiveness of each module by comparing with the following variants step by step: i) linear: Use linear layer to align feature dimension only; ii) attn: Based on i, further use a single self-attention block to align the distribution; iii) cross x4: we instead use learnable query and 4 cross-attention layers to learn the MAE prior. ASUKA adopts 4 self-attention blocks. Results are shown in Tab. 2(b). The self-attention block shows improved results compared with only align dimension and cross-attention block. Using 4 self-attention blocks improves the capacity of the alignment module.
Ablation of decoder. For the decoder, we compare ASUKA with i) VAE: the decoder used in SD; ii) + cond.: the decoder conditioned on unmasked image [62]; iii) + color: only use color augmentation to train the decoder; Results are in Tab. 2(d), showing the superiority of our augmented decoder.
Table 2: Ablation stuides of ASUKA.
(a) Comparison of ASUKA using pre-trained (p.t.) MAE v.s. fine-tuned (f.t.) MAE.
@@ 5.2 Further Analysis of ASUKA
p.t. 0.151 11.513 0.354 0.258
f.t. 0.150 11.460 0.368 0.256
(b) Ablation of different alignment modules. Linear adopts linear layer; attn adopts a self-attention layer; cross x4 adopts 4 cross-attention layers; ASUKA adopts 4 self-attention layers.
@@
linear 0.155 11.934 0.361 0.227
attn 0.152 11.613 0.362 0.234
cross x4 0.152 11.762 0.368 0.238
ASUKA 0.150 11.460 0.368 0.256
(c) User-study of top-1 ratio among all the inpainting results. Context-stability (C.S.) measures the coherence between masked region and unmasked surroundings, with a preference of not generating new elements; Visual-consistency (V.C.) measures the color consistency.
Model C.S.(%) V.C.(%)
@@
ASUKA 39.43 40.05
(d) Comparison of different decoders for SD. VAE [39] is the original decoder used by SD; + cond. [62] is the decoder conditioned on unmasked image; + color uses the color augmentation strategy to perform local harmonization task; Ours further combines latent augmentation strategy to handle the gap between generated latent and real latent.
[…]
VAE 0.156 11.949 0.343 0.208
+ cond. 0.151 11.634 0.361 0.231
+ color 0.152 11.603 0.357 0.229
Ours 0.150 11.460 0.368 0.256
User-Study. To evaluate the user preference on inpainting algorithms, we conduct an user-study. Specifically, we randomly select 40 testing images. We ask the user to select the best inpainting results from the following perspectives respectively: i) Context-stability (C.S.): the generated region should be context-stable with surrounding unmasked region, with a preference of not generating new elements; ii) Visual-consistency (V.C.) : the visual consistency between masked and unmasked regions. We collect 100 valid anonymous questionnaire results, and report the average selection ratio among all the inpainting algorithms in Tab. 2(c). This result validate the efficacy of ASUKA on alignment with human preference.
Limitation: The "curse" of self-attention. The primary limitation of ASUKA stems from the inefficacy of the MAE prior, mainly due to issues within the self-attention module. Specifically, as shown in Fig. 9, the presence of multiple similar objects in an image may lead the MAE to incorrectly predict a similar object in the masked region, conflicting with the goal of object removal. Notably, this curse of self-attention significantly impacts diffusion-based generative models. It results in the inability to accurately follow "blank paper" text prompts, even when employing a substantial classifier-free guidance scale of 9. This issue is not unique to SD but is also a common problem in other advanced text-guided diffusion models, such as OpenAI’s DALL-E 2 [38] and Adobe’s FileFly [1]. Nevertheless, ASUKA has the potential to circumvent this issue by modifying the MAE prior, for instance, by instead using a blank paper image as the input to MAE prior. A more comprehensive solution would involve extra control on self-attention layers in diffusion models, which we leave as future work.
6 Conclusion
In this paper, we proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) to achieve context-stable and visual-consistent inpainting via frozen Stable Diffusion inpainting model (SD). To achieve context-stability, we adopt a reconstruction-based masked auto-encoder (MAE) as the context-stable prior for masked region purely from unmasked region. Then we align the context-stable prior to frozen SD with the proposed alignment module. To achieve visual-consistency, we resolve the mask-unmask color inconsistency of SD in the decoding process. We train an unmask-region conditioned KL-VAE decoder to perform local harmonization during the decoding process. To validate the efficacy of inpainting algorithms in different image domains and mask types, we introduce an evaluation dataset, named as MISATO, from existing datasets. ASUKA enjoys context-stable and visual-consistent inpainting results and superior than state-of-the-art inpainting algorithms and SD-based variants.

### v3

@@ 7 Details about MISATO
Combined together, MISATO contains 2000 examples from four inpainting domains, indoor, outdoor landscape, building, and background, as shown in Fig. 10. we adopt the masking strategy as in Sec. 3.1.1 but excluding the rectangle and complement rectangle masks. The masking ratio is set as $[0.2,0.8]$.
Figure 10: Different image domains in MISATO.
8 Implementation Details
We use Places2 [101] to train ASUKA. For the MAE [31] used in ASUKA, we train on images of size $256^{2}$, which is efficient and produce context-stable guidance for generative models to generate high-resolution images. We fine-tune the MAE with a batch size of 1024. We train the alignment module with AdamW [53] of learning rate 5e-2 with the standard diffusion objective. We set $p$ as $100\%$ and linearly decay it to $10\%$ in the first 2K training steps and then freeze. For SD’s decoder, we fine-tune from [103] for 50K steps with a batch size of 40 and learning rate of 8e-5 with cosine decay. For FLUX’s decoder, we fine-tune from the original decoder with the same setup. We use ColorJitter for color augmentation, with brightness 0.15, contrast 0.2, saturation 0.1, and hue 0.03.
9 Further Analysis
Table 5: Comparison of ASUKA with text-guided SD
@@ 9 Further Analysis
SD (BLIP2) 0.163 12.536 0.370 0.225 0.880 70.846
ASUKA-SD 0.150 11.495 0.423 0.312 0.958 47.753
Comparison with text-guided inpainting
We compare ASUKA with text-guided SD model, as shown in Tab. 5. We run SD inpainting sing text captions generated by BLIP2 [43]. ASUKA performs better, since captions describe the entire image, while MAE focuses on reconstructing only the masked region, leading to more precise guidance.
Table 6: Ablation of $p$
@@
p=0 0.155 11.804 0.403 0.288 0.940 48.032
p=1 0.152 11.734 0.394 0.296 0.947 47.997
linear decay p 0.152 11.558 0.405 0.307 0.955 47.814
Ours 0.150 11.495 0.423 0.312 0.958 47.753
Ablation of $p$
We analyze how different values of $p$ affect ASUKA in Tab. 6. The results show that our warm-up and freeze strategy outperforms other approaches.
Table 7: Additional results on benchmark datasets
@@
CelebA-HQ SD 0.132 11.968 0.282 0.101 0.939 42.870
ASUKA-SD 0.129 10.190 0.293 0.134 0.941 40.503
FFHQ SD 0.139 2.235 0.371 0.197 0.944 43.529
ASUKA-SD 0.131 2.060 0.386 0.205 0.955 30.848
Additional Results
We further compare ASUKA with standard SD on two additional datasets: CelebA-HQ [36] and FFHQ [37]. As shown in Tab. 7, these results provide more evidence of ASUKA’s effectiveness.
Table 8: Our Decoder in Text-Guided Inpainting.
[…]
SD 0.297 0.180 30.255 0.312 0.930 57.136
ASUKA-SD 0.298 0.175 29.350 0.350 0.931 38.123
Our Decoder in Text-Guided Inpainting
To test the generalizability of our decoder, we evaluate it on text-guided inpainting tasks. We compare our decoder with the original SD decoder using 1,000 randomly sampled images from “jackyhate/text-to-image-2M” [105]. The results in Tab. 8 confirm its effectiveness for general inpainting tasks.
Table 9: Effect of each module.
[…]
SD w/ MAE 0.157 12.093 0.397 0.236 0.953 62.845
SD w/ decoder 0.159 12.075 0.411 0.283 0.954 49.376
ASUKA-SD 0.150 11.495 0.423 0.312 0.958 47.753
Ablation on independent modules
To understand the contribution of each module in ASUKA, we evaluate SD with the proposed modules added separately. The results, shown in Tab. 9, highlight the effectiveness of each module.
Table 10: Comparison of ASUKA using pre-trained MAE v.s. fine-tuned MAE.
[…]
pre-trained 0.151 11.513 0.354 0.258
fine-tuned 0.150 11.460 0.368 0.256
Ablation of MAE prior
We compare our fine-tuned MAE with directly adopting the MAE trained in [10]. To this end, we train ASUKA with the MAE in [10] using the same training strategy and compare the results in Tab. 10. Results suggest the improvements of fine-tuning MAE, especially on FID and U-IDS. This improvement comes from the better adaptation on the real-world masks.
Table 11: User-study of top-1 ratio among all the inpainting results.
Model UOM (%) CC(%)
@@
ASUKA-SD 39.43 40.05
User-study
To evaluate the user preference on inpainting algorithms, we conduct an user-study. Specifically, we randomly select 40 testing images. We ask the user to select the best inpainting results from the following perspectives respectively: i) Unwanted-object-mitigation (UOM): the generated region should be context-stable with surrounding unmasked region, with a preference of not generating new elements; ii) Color-consistency (CC) : the color consistency between masked and unmasked regions. We collect 100 valid anonymous questionnaire results, and report the average selection ratio among all the inpainting algorithms in Tab. 11. This result validate the efficacy of ASUKA on alignment with human preference.
Limitation: The "curse" of self-attention
The primary limitation of ASUKA stems from the inefficacy of the MAE prior, mainly due to issues within the self-attention module. Specifically, as shown in Fig. 11, the presence of multiple similar objects in an image may lead the MAE to incorrectly predict a similar object in the masked region, conflicting with the goal of object removal. Notably, this curse of self-attention significantly impacts diffusion-based generative models. It results in the inability to accurately follow "blank paper" text prompts, even when employing a substantial classifier-free guidance scale of 9. This issue is not unique to SD but is also a common problem in other advanced text-guided diffusion models, such as OpenAI’s DALL-E 2 [65] and Adobe’s FileFly [18]. Nevertheless, ASUKA has the potential to circumvent this issue by modifying the MAE prior, for instance, by instead using a blank paper image as the input to MAE prior. A more comprehensive solution would involve extra control on self-attention layers in diffusion models, which we leave as future work.
Potential negative impact
As an image editing tool, our proposed ASUKA will generate images based on user intentions for masking specific parts of the image, potentially resulting in unrealistic renderings and posing a risk of misuse.